### v2

@@ Abstract
Context:The three-body problem (3BP) poses a longstanding challenge in physics and celestial mechanics. Despite the impossibility of obtaining general analytical solutions, statistical theories have been developed based on the ergodic principle. This assumption is justified by chaos, which is expected to fully mix the accessible phase space of the 3BP.
Aims:This study probes the presence of regular (i.e. non chaotic) trajectories within the 3BP and assesses their impact on statistical escape theories.
Methods:Using numerical simulations, we establish criteria for identifying regular trajectories and analyse their impact on statistical outcomes.
Results:Our analysis reveals that regular trajectories occupy up to 32% of the phase space, and their outcomes defy the predictions of statistical escape theories. The coexistence of regular and chaotic regions at all scales is characterized by a multi-fractal behaviour. Integration errors manifest as numerical chaos, artificially enhancing the mixing of the phase space and affecting the reliability of individual simulations, yet preserving the statistical correctness of an ensemble of realizations.
Conclusions:Our findings underscore the challenges in applying statistical escape theories to astrophysical problems, as they may bias results by excluding the outcome of regular trajectories. This is particularly important in the context of formation scenarios of gravitational wave mergers, where biased estimates of binary eccentricity can significantly impact estimates of coalescence efficiency and detectable eccentricity.
Isles of regularity in a sea of chaos amid the gravitational three-body problem
@@
No matter how complex a three-body interaction may appear, its evolution can be subdivided into two distinct states that the system alternates between (see Figure 1). The first state occurs when all three particles interact within a confined region of space, devoid of any fixed hierarchy among them. We refer to this state of the interaction as a ‘democratic resonance’, as the particles freely exchange energy and angular momentum among each other (see Heggie & Hut, 1993). Democratic resonances cannot persist indefinitely, because this configuration is inherently unstable, and the system will eventually form a temporary hierarchy, transitioning into the second state, which we term ‘excursion’. Excursions occur when two bodies form a temporary binary and the third body recoils while remaining bound to the binary centre of mass. In this state, the system forms an unstable, temporary hierarchical triple, which will revert back to the democratic resonance state once the single body returns to interact with the binary. A three-body interaction cycles between these two states until the system finally breaks up into an unbound binary-single pair. The breakup can only occur from the democratic resonance state, because during an excursion, the system can be considered as non-interacting. The physical origin of chaos in the non-hierarchical 3BP arises from the close gravitational interactions during the democratic resonances.
@@
Fortunately, it is possible to leverage chaos to find statistical solutions of the 3BP, also called statistical escape theories, which have origin in the pioneering work of Monaghan (1976a). His approach, which is based on the microcanonical ensemble of statistical mechanics, laid the foundation for an exploration of the interplay between deterministic chaos and probabilistic dynamics in celestial mechanics.
@@
The key idea of statistical escape theories is that chaotic interactions mix the phase space of the 3BP, so that over the lifetime of the system, every available state will be eventually visited, filling the phase space in a uniform and random sense. Except conserved quantities (energy and angular momentum vector), the system has lost memory of its initial conditions due to chaos and, at any given time, it can be found at any point in the accessible phase space. This is the definition of an ergodic system in statistical mechanics, meaning that the time average of some property over the trajectory is equal to the ensemble average over the phase space volume (Birkhoff, 1931). In the case of the non-hierarchical 3BP, it means that the probability of outcome properties are proportional to the phase space that is accessible at the breakup time of the triple system. Given energy and angular momentum vector, it is then possible to characterize and sample the accessible phase space at the moment of breakup to obtain the distributions of final outcome properties (Monaghan, 1976a, b).
@@
Statistical escape theories typically provide the closed-form expression for the tri-variate distribution of outcome properties , expressed in terms of final binary binding energy , angular momentum and the angle between the total angular momentum and the angular momentum of the ejected single . The original statistical escape theory by Monaghan (1976a) has been improved and refined over the years, yielding a remarkably good agreement with the results of numerical experiments (Standish, 1972a; Saslaw et al., 1974; Valtonen, 1974; Monaghan, 1976b; Valtonen et al., 2005; Stone & Leigh, 2019; Manwadkar et al., 2020, 2021; Ginat & Perets, 2021).
@@
However, statistical escape theories suffer from the assumption that the phase space of the 3BP is fully mixed. It has been highlighted that the 3BP is not fully chaotic, and its phase space is divided between chaotic and regular (i.e. non-chaotic) trajectories (Shevchenko, 1998a, b, 2010; Manwadkar et al., 2020; Parischewsky et al., 2023). Therefore, we can expect that the predictions of the statistical escape theories clash with the outcomes of regular 3BP interactions. Only recently, a new statistical theory of the 3BP, not based on ergodicity, has been proposed with the aim of solving this issue (Kol, 2021). Even though the theory is not complete yet, preliminary results show an improved agreement of the predicted ejection probabilities with the numerical experiments, compared to previous theories (Manwadkar et al., 2021, 2023).
@@
In this paper, we explore the presence of regular trajectories in the 3BP and its impact on the statistical predictions of the 3BP escape theories. Section 2 details the setup of our simulations and our numerical model. Criteria for identifying regular trajectories in the initial phase space are outlined in Section 3, which also illustrates the influence of regular regions on the statistical outcomes of the 3BP. Section 4 delves into the fine-grained mixing between regular and chaotic regions occurring at all scales, characterizing the multi-fractal nature of this mixing. Integration errors and their role in spurious phase space mixing, referred to as numerical chaos, are examined in Section 5. Section 6 highlights the extreme sensitivity of the 3BP to perturbations below the Planck length, where mixing between regularity and chaos persists. In Section 7 we highlight the astrophysical implications of our work. Finally, our conclusion and outlooks are summarized in Section 8.
@@ 2 Numerical methods
We perform a suite of numerical simulations of three-body interactions. The interactions begin with a binary with a fixed semi-major axis of and a single, which undergo an interaction with a given total energy, angular momentum and particle masses. The initial setup is illustrated in Figure 2. The center of mass of the binary and the single are at rest in the center-of-mass frame of the system, at a distance of .
@@
We sample the initial parameter space uniformly in the - space, as shown in Figure 3. We simulate realizations for the whole parameter space (, ). We also perform a series of 14 zoom-ins into the - space, by sampling initial conditions in 1/25th of the area of the previous zoom level (i.e., a shrinking factor of 5 in each dimension). The sequence of zoom-ins is displayed in Figure 5 and Figure 6.
We integrate the motion of the 3 particles assuming no particle size and no collision, by means of the tsunami integrator (Trani et al., 2019; Trani & Spera, 2023). tsunami uses a combination of algorithmic techniques (Stoer & Bulirsch, 1980; Mikkola & Aarseth, 1993; Mikkola & Tanikawa, 1999) to improve the accuracy and speed of the -body integration, which results in total energy errors .
@@ 3 Regular trajectories elude statistical escape theories
The top panel of Figure 3 shows the initial phase space map, color-coded by the identity of the escaping particle. In this space, neighbouring points share similar initial conditions. If the 3BP was fully ergodic, we should see a mix of colors everywhere, but instead we observe four large regions of uniform colors, two big ones at and , and two small ones at and , that we term regular islands. These regular islands are constituted by regular trajectories (as opposed to chaotic trajectories), where the final state of the interaction smoothly maps onto the initial conditions. These correspond to escapes of either of the two members of the binary. The blue and green regular islands present a translational quasi-symmetry by 180 along the -axis, which correspond to the binary phase swapping the two binary members. The slightly unequal masses prevent it from being a perfect symmetry, and make the small green island (escape of the most massive particle) larger than the small blue island, and the big blue island more lopsided at high than the green one.
To quantify how much area is occupied by the regular regions, we employ a nearest neighbour classification scheme. For each point in the phase space, we calculate the fraction of neighbours that have the same color as the point in consideration. The bottom panel of Figure 3 color-codes the realizations based on for .ly This criterion appears to effectively identify the major regular islands while also revealing numerous narrow, fine-grained regular regions outside these islands, which we refer to as regular stripes.
If we define the regular trajectories as those that have , we find that about 32% of the phase space is constituted by regular regions, with the blue () and green () regular regions constituting about 10% and 13% of the phase space each, and the remaining 9% constituted by escapes of the lighter body (red, ). The proportion of regular trajectories increases with the mass of the particles. This is intuitively expected, as for , the motion of the binary becomes a Keplerian orbit (and thus regular by definition), while the motion of the point mass remains chaotic.
@@
An unexplored aspect thus far is how these regular trajectories influence the distribution of outcome properties. If we were to color-code the outcome properties such as the final binary semi-major axis in the initial phase space map, smooth transitions would appear corresponding to the regular regions. It is therefore not surprising that such regular regions have a considerable impact on the outcome properties, as shown in Figure 4, which compares the marginalized distributions of semi-major axis and eccentricity to the theoretical predictions of Stone & Leigh (2019).
The distributions of final outcomes, when considering all realizations, are at odds with the theoretical expectations for the ejected masses of and (blue and green lines). The discrepancy with the theory is particularly visible in the distributions of final eccentricities, which are more peaked at . Additionally, the semi-major axes distributions are skewed to large values, and clearly deviate from the simple power-law behaviour that we expect from the theory. However, the outcome distribution of the low-mass escapes (red lines) broadly agrees with theoretical expectations. We attribute this agreement to the fact that only 22% of all the red escape trajectories are regular, despite their fraction over total being 9%. In contrast, regular trajectories account for about 34% and 48% of the trajectories for the green and blue escapes, respectively. In other words, not only does the total fraction of regular interactions increase for increasing mass, but also the relative fraction for each escaper mass.
@@
Our findings have significant implications in the astrophysical application of statistical escape theories. Astrophysical predictions based on statistical theories will be inherently biased, because they exclude the outcomes of regular interactions. This issue is particularly concerning in the context of formation scenarios of gravitational wave mergers (e.g., Kritos et al., 2022; Mapelli et al., 2022; Ginat & Perets, 2023), where the final binary eccentricity, known to dramatically affect the gravitational wave radiation mechanism (at high , the coalescence time , see Peters, 1964), is significantly different in the regular regions of the phase space. We delve deeper into this aspect in Section 7.
@@ 4 Chasing Laplace’s demon
In addition to the four big regular islands, Figure 3 highlights the presence of finer structures that look like narrow stripes. Between these structures, are there regions without regular features, where all trajectories have outcomes uncorrelated with their neighbours? In Figure 5 we show the initial phase space maps for the sequence of zoom-ins, beginning from a region of apparent chaos. As we zoom in, we observe the emergence of regular structures from what seemed at first glance to be a chaotic region. The regular structures are self-similar, taking the appearance of slanted stripes.
Similar patterns are visible if we color code by the lifetime of the system, as in Figure 6. Uniform regions in the ejected color map correspond to regions where the lifetime of the system is extremely short, comparable to the free fall time of the binary-single system . Regular interactions have very short lifetimes because the system breaks up after a single close encounter between all three-particles, which is not enough to make the system highly chaotic. Several studies have indeed found that there is an approximate power-law relation between the Lyapunov time and the lifetime of the gravitational 3BP (Lecar et al., 1992; Orlov et al., 2010; Mikkola & Tanikawa, 2007; Urminsky & Heggie, 2009). The main reason is that a triple system with a longer lifetime has more opportunities to undergo democratic resonances than a short-lived one. On the other hand, a long-lived system may also have spent most of its time in an excursion state, so the relation between and is only approximate (Manwadkar et al., 2020).
@@
First, from Figure 7 we observe that the fraction of regular interactions , as measured from our -neighbours scheme, changes depending on which zoom level we are considering. Overall, the total percentage of regular interactions varies between 32% and 15% across the zoom levels. The relative fraction of regular trajectories for each escaper mass changes significantly over the zoom range. The low-mass escapes, initially having the lowest relative fraction of regular trajectories in the outer box, display the highest relative fraction in all zoom levels except the 9th one. This discrepancy is evident from Figure 5, where the 9th zoom-in prominently features a large green stripe, constituting the majority of the regular space. Conversely, the high-mass escapes, initially exhibiting the highest fraction of regular trajectories in the outer box, display the least relative fraction in most of the zoom levels. This is primarily due to the paucity of high-mass escapes in the zoom-ins, compounded by our numerical resolution limitations as detailed in Section 5. The varying fraction of regular interactions from the outer box to the zoom-ins underscores the qualitative difference in the system behaviour across different scales.
@@
Therefore, we adopt instead the 2-point correlation dimension, which is known to be accurate even when only a small number of points is available (Grassberger & Procaccia, 1983). This technique works by counting the number of pairs whose distance is less then . Then, for small , grows as a power-law, i.e. , where is the two-point correlation dimension, which is guaranteed to be close (strictly speaking, equal to or less than) the fractal dimension of the set.
@@
As shown in the left-panel of Figure 8, measuring for all the points in a box in phase space yields a dimension close to the expected one, i.e. . We attribute the slightly smaller value we obtained () to numerical artefacts in the pseudo-random generations of the realizations in the box, which produce some small degree of clustering that is expected to occur using the Mersenne-Twister pseudo-random generator. Note how the curve deviates from a power-law at the extrema, where we either approach the resolution of our set or the large-scale boundary of the box. To prevent these boundaries from affecting our least squared fit to the power-law exponent, we weight the data using a Gaussian function centered on the mean scale.
@@
We then select only the regular set, here again defined as the ensemble of realizations with a same-colour fraction among the nearest neighbours, to select only the most uniform region. We find a fractal dimension that is smaller than 2 at all zoom levels, which points out the fractal nature of regular trajectories in phase space, because it is fractional and smaller than the dimensions of the embedding space. However, each zoom-level presents a different fractal index, as seen in the right panel of Figure 8. Overall, the fractal dimension changes across the zoom-ins, which confirms our hypothesis of the multi-fractal nature of chaos in the 3BP. Interestingly, remains constant across the 3rd, 4th, and 5th zoom-ins, which also exhibit remarkably similar features in Figure 6 (thick regular stripes). Similarly, the 6th, 7th, and 8th zoom-ins, characterized by thin regular stripes, also display comparable fractal dimensions. The most significant change in occurs with the 9th zoom-in, featuring a large regular stripe, resulting in higher values, indicative of a more two-dimensional structure.

### v3

@@ Abstract
Context. The three-body problem (3BP) poses a longstanding challenge in physics and celestial mechanics. Despite the impossibility of obtaining general analytical solutions, statistical theories have been developed based on the ergodic principle. This assumption is justified by chaos, which is expected to fully mix the accessible phase space of the 3BP.
Aims. This study probes the presence of regular (i.e. non-chaotic) trajectories within the 3BP and assesses their impact on statistical escape theories.
Methods. Using three-body simulations performed with the accurate, regularized code tsunami, we established criteria for identifying regular trajectories and analysed their impact on statistical outcomes.
Results. Our analysis reveals that regular trajectories occupy a significant fraction of the phase space, ranging from 28% to 84% depending on the initial setup, and their outcomes defy the predictions of statistical escape theories. The coexistence of regular and chaotic regions at all scales is characterized by a multi-fractal behaviour. Integration errors manifest as numerical chaos, artificially enhancing the mixing of the phase space and affecting the reliability of individual simulations, yet preserving the statistical correctness of an ensemble of realizations.
Conclusions. Our findings underscore the challenges in applying statistical escape theories to astrophysical problems, as they may bias results by excluding the outcome of regular trajectories. This is particularly important in the context of formation scenarios of gravitational wave mergers, where biased estimates of binary eccentricity can significantly impact estimates of coalescence efficiency and detectable eccentricity.
@@
No matter how complex a three-body interaction may appear, its evolution can be subdivided into two distinct states between which the system alternates (see Figure 1). The first state occurs when all three particles interact within a confined region of space, devoid of any fixed hierarchy among them. We refer to this state of the interaction as a ‘democratic resonance’, as the particles freely exchange energy and angular momentum among each other (see Heggie & Hut, 1993). Democratic resonances cannot persist indefinitely, because this configuration is inherently unstable, and the system eventually forms a temporary hierarchy, transitioning into the second state, which we term ‘excursion’. Excursions occur when two bodies form a temporary binary and the third body recoils while remaining bound to the binary centre of mass. In this state, the system forms an unstable, temporary hierarchical triple, which reverts back to the democratic resonance state once the single body returns and interacts with the binary. A three-body interaction cycles between these two states until the system finally breaks up into an unbound binary-single pair. The breakup can only occur from the democratic resonance state, because during an excursion, the system can be considered as non-interacting. The physical origin of chaos in the non-hierarchical 3BP arises from the close gravitational interactions during the democratic resonances.
@@
Fortunately, it is possible to leverage chaos to find statistical solutions for the 3BP, also called statistical escape theories, which have an origin in the pioneering work of Monaghan (1976a). His approach, which is based on the microcanonical ensemble of statistical mechanics, laid the foundation to explore the interplay between deterministic chaos and probabilistic dynamics in celestial mechanics.
@@
The key idea of statistical escape theories is that chaotic interactions mix the phase space of the 3BP, so that over the lifetime of the system, every available state is eventually visited, filling the phase space in a uniform and random sense. Except for conserved quantities (the energy and angular momentum vector), the system has lost memory of its initial conditions due to chaos and, at any given time, it can be found at any point in the accessible phase space. This is the definition of an ergodic system in statistical mechanics, meaning that the time average of some property over the trajectory is equal to the ensemble average over the phase-space volume (Birkhoff, 1931). In the case of the non-hierarchical 3BP, this means that the probability of the outcome properties are proportional to the phase space that is accessible at the breakup time of the triple system. Given the energy and angular momentum vector, it is then possible to characterize and sample the accessible phase space at the moment of breakup to obtain the distributions of the final outcome properties (Monaghan, 1976a, b).
@@
Statistical escape theories typically provide the closed-form expression for the tri-variate distribution of the outcome properties, expressed in terms of final binary binding energy , angular momentum , and the angle between the total angular momentum and the angular momentum of the ejected single . The original statistical escape theory by Monaghan (1976a) has been improved and refined over the years, yielding a remarkably good agreement with the results of numerical experiments (Standish, 1972a; Saslaw et al., 1974; Valtonen, 1974; Monaghan, 1976b; Valtonen et al., 2005; Stone & Leigh, 2019; Manwadkar et al., 2020, 2021; Ginat & Perets, 2021).
@@
However, statistical escape theories suffer from the assumption that the phase space of the 3BP is fully mixed. It has been highlighted that the 3BP is not fully chaotic, and its phase space is divided between chaotic and regular (i.e. non-chaotic) trajectories (Shevchenko, 1998a, b, 2010; Manwadkar et al., 2020; Parischewsky et al., 2023). Therefore, we can expect that the predictions of the statistical escape theories clash with the outcomes of regular 3BP interactions. Only recently, a new statistical theory of the 3BP, not based on the microcanonical ensemble, has been proposed with the aim of solving this issue (Kol, 2021). Simulations show an improved agreement of the predicted ejection probabilities with the numerical experiments, compared to previous theories (Manwadkar et al., 2021, 2023).
@@
In this paper, we explore the presence of regular trajectories in the 3BP and its impact on the statistical predictions of the 3BP escape theories. Section 2 details the setup of our simulations and our numerical model. Criteria for identifying regular trajectories in the initial phase space are outlined in Section 3, which also illustrates the influence of regular regions on the statistical outcomes of the 3BP. Section 4 delves into the fine-grained mixing between regular and chaotic regions occurring at all scales, characterizing the multi-fractal nature of this mixing. Integration errors and their role in spurious phase-space mixing, referred to as numerical chaos, are examined in Section 5. Section 6 highlights the extreme sensitivity of the 3BP to perturbations below the Planck length, where mixing between regularity and chaos persists. In Section 7 we highlight the astrophysical implications of our work. Finally, our conclusion and outlooks are summarized in Section 8.
@@ 2 Numerical methods
We perform a suite of numerical simulations of three-body interactions. The interactions begin with a binary with a fixed semi-major axis of and a single, which undergo an interaction with a given total energy, angular momentum and particle masses. The initial setup is illustrated in Figure 2. The centre of mass of the binary and the single are at rest in the centre-of-mass frame of the system, at a distance of .
@@
We sample the initial parameter space uniformly in the - space, as shown in Figure 3. We simulate realizations for the whole parameter space (, ). We also perform a series of 14 zoom-ins into the - space, by sampling initial conditions in 1/25th of the area of the previous zoom level (i.e. a shrinking factor of 5 in each dimension). Both and are sampled uniformly. However, when calculating the statistical properties of the sets, such as the fraction of regular interactions, we weight each realization by to ensure that our statistics are consistent with being uniform in .
We integrate the motion of the three particles assuming no particle size and no collision, by means of the tsunami integrator (Trani et al., 2019; Trani & Spera, 2023). tsunami uses a combination of algorithmic techniques (Stoer & Bulirsch, 1980; Mikkola & Aarseth, 1993; Mikkola & Tanikawa, 1999) to improve the accuracy and speed of the -body integration, which results in total energy errors .
@@ 3 Regular trajectories elude statistical escape theories
The top panel of Figure 3 shows the initial phase space map, colour-coded by the identity of the escaping particle. In this space, neighbouring points share similar initial conditions. If the 3BP was fully ergodic, we should see a mix of colours everywhere, but instead we observe four large regions of uniform colours, two large ones at and , and two small ones at and , that we term regular islands. These regular islands are constituted by regular trajectories (as opposed to chaotic trajectories), where the final state of the interaction smoothly maps onto the initial conditions. These correspond to escapes of either of the two members of the binary. The blue and green regular islands present a translational quasi-symmetry by 180∘ along the -axis, which correspond to the binary phase swapping the two binary members. The slightly unequal masses prevent it from being a perfect symmetry, and make the small green island (escape of the most massive particle) larger than the small blue island, and the large blue island more lopsided at high than the green one.
To quantify how much area is occupied by the regular regions, we employ a nearest neighbour classification scheme. For each point in the phase space, we calculate the fraction of neighbours that have the same colour as the point in consideration. The bottom panel of Figure 3 colour-codes the realizations based on for . This criterion appears to effectively identify the major regular islands while also revealing numerous narrow, fine-grained regular regions outside these islands, which we refer to as regular stripes.
If we define the regular trajectories as those that have , we find that about 37% of the phase space is constituted by regular regions, with the blue () and green () regular regions constituting about 14% and 13% of the phase space each, and the remaining 9% constituted by escapes of the lighter body (red, ). The proportion of regular trajectories increases with the mass of the particles. This is intuitively expected, as for , the motion of the binary becomes a Keplerian orbit (and thus regular by definition), while the motion of the point mass remains chaotic.
@@
An unexplored aspect thus far is how these regular trajectories influence the distribution of outcome properties. If we were to colour-code the outcome properties such as the final binary semi-major axis in the initial phase space map, smooth transitions would appear corresponding to the regular regions. It is therefore not surprising that such regular regions have a considerable impact on the outcome properties, as shown in Figure 4, which compares the marginalized distributions of semi-major axis and eccentricity to the theoretical predictions of Stone & Leigh (2019).
The distributions of final outcomes, when considering all realizations, are at odds with the theoretical expectations for the ejected masses of and (blue and green lines). The discrepancy with the theory is particularly visible in the distributions of final eccentricities, which are more peaked at . Additionally, the semi-major axes distributions are skewed to large values, and clearly deviate from the simple power-law behaviour that we expect from the theory. However, the outcome distribution of the low-mass escapes (red lines) broadly agrees with theoretical expectations. We attribute this agreement to the fact that only 22% of all the red escape trajectories are regular, despite their fraction over the total being 9%. In contrast, regular trajectories account for about 42% and 52% of the trajectories for the green and blue escapes, respectively. In other words, not only does the total fraction of regular interactions increase for increasing mass, but also the relative fraction for each escaper mass.
@@
Our findings have significant implications in the astrophysical application of statistical escape theories. Astrophysical predictions based on statistical theories will be inherently biased, because they exclude the outcomes of regular interactions. This issue is particularly concerning in the context of formation scenarios of gravitational wave mergers (e.g. Kritos et al., 2022; Mapelli et al., 2022; Ginat & Perets, 2023), where the final binary eccentricity, known to dramatically affect the gravitational wave radiation mechanism (at high , the coalescence time , see Peters, 1964), is significantly different in the regular regions of the phase space. We delve deeper into this aspect in Section 7.
@@ 4 Chasing Laplace’s demon
In addition to the four large regular islands, Figure 3 highlights the presence of finer structures that look like narrow stripes. We explored whether there are regions between these structures without regular features, where all trajectories have outcomes uncorrelated with their neighbours. In Figure 5 we show the initial phase space maps for the sequence of zoom-ins, beginning from a region of apparent chaos. As we zoom in, we observe the emergence of regular structures from what seemed at first glance to be a chaotic region. The regular structures are self-similar, taking the appearance of slanted stripes.
Similar patterns are visible if we colour code by the lifetime of the system, as in Figure 6. Uniform regions in the ejected colour map correspond to regions where the lifetime of the system is extremely short, comparable to the free fall time of the binary-single system . Regular interactions have very short lifetimes because the system breaks up after a single close encounter between all three-particles, which is not enough to make the system highly chaotic. Several studies have indeed found that there is an approximate power-law relation between the Lyapunov time and the lifetime of the gravitational 3BP (Lecar et al., 1992; Orlov et al., 2010; Mikkola & Tanikawa, 2007; Urminsky & Heggie, 2009). The main reason is that a triple system with a longer lifetime has more opportunities to undergo democratic resonances than a short-lived one. On the other hand, a long-lived system may also have spent most of its time in an excursion state, so the relation between and is only approximate (Manwadkar et al., 2020).
@@
First, from Figure 7 we observe that the fraction of regular interactions , as measured from our -neighbours scheme, changes depending on which zoom level we are considering. Overall, the total percentage of regular interactions varies between 37% and 15% across the zoom levels. The relative fraction of regular trajectories for each escaper mass changes significantly over the zoom range. The low-mass escapes, initially having the lowest relative fraction of regular trajectories in the outer box, display the highest relative fraction in all zoom levels except the 9th one. This discrepancy is evident from Figure 5, where the 9th zoom-in prominently features a large green stripe, constituting the majority of the regular space. Conversely, the high-mass escapes, initially exhibiting the highest fraction of regular trajectories in the outer box, display the least relative fraction in most of the zoom levels. This is primarily due to the paucity of high-mass escapes in the zoom-ins, compounded by our numerical resolution limitations as detailed in Section 5. The varying fraction of regular interactions from the outer box to the zoom-ins underscores the qualitative difference in the system behaviour across different scales.
@@
Therefore, we adopt instead the 2-point correlation dimension, which is known to be accurate even when only a small number of points is available (Grassberger & Procaccia, 1983). This technique works by counting the number of pairs whose distance is less then . Then, for small , grows as a power law, that is , where is the two-point correlation dimension, which is guaranteed to be close (strictly speaking, equal to or less than) the fractal dimension of the set. In a finite box, points near the edges have fewer neighbours within a distance compared to points well inside the box. This reduction in neighbours leads to an underestimation of near the edges. To address this, we applied an edge-correction factor to , estimated by comparing from a uniform distribution with and without a toroidal topology at the boundaries.
The fractal dimension is a measure of the complexity of a fractal surface embedded in a space of dimension , indicating how it scales with size. It has found applications in several astrophysical scenarios, for instance, to measure the clustering degree in clusters of proto-stars (Larson, 1995; Cartwright & Whitworth, 2004; Ballone et al., 2021). In our case, the initial phase space of our 3BP setup resides in . A fractal dimension of implies that the considered set is not fractal but smooth. This suggests that the underlying dynamics is regular rather than chaotic, as the set uniformly populates the embedding space. This is clear when examining the large islands in Figure 3, which are caused by the presence of a smooth mapping between initial and final phase spaces, a characteristic of regular dynamics. The lower the value of , the sparser the set becomes, meaning it fills less of the embedding space. It is also possible that the correlation function may not be adequately described by a single power-law, suggesting that the fractal exhibits varying levels of complexity or self-similarity across different scales. In such instances, the fractal is more accurately characterized by a spectrum of fractal dimensions, giving rise to what is known as a multi-fractal.
@@
As shown in the left-panel of Figure 8, measuring for uniformly distributed points in a box yields a dimension consistent with the expected one, that is . It is important to note how the curve deviates from a power-law at the extrema, where we either approach the resolution of our set or the large-scale boundary of the box. To prevent these boundaries from affecting our least squared fit to the power-law exponent, we weight the data using a Gaussian function centred on the mean scale.
@@
We then select only the regular set, here again defined as the ensemble of realizations with a same-colour fraction among the nearest neighbours, to select only the most uniform region. We find a fractal dimension that is smaller than 2 at all zoom levels, which points out the fractal nature of regular trajectories in the phase space, because it is fractional and smaller than the dimensions of the embedding space. However, each zoom-level presents a different fractal index, as seen in the right panel of Figure 8. Overall, the fractal dimension changes across the zoom-ins, which confirms our hypothesis of the multi-fractal nature of chaos in the 3BP. Interestingly, remains constant across the 3rd, 4th, and 5th zoom-ins, which also exhibit remarkably similar features in Figure 6 (thick regular stripes). Similarly, the 6th, 7th, and 8th zoom-ins, characterized by thin regular stripes, also display comparable fractal dimensions. The fractal dimension for zoom levels 6,7,8 is smaller than that of zoom levels 3,4,5, which are consistent with having sparser, thinner stripes. The most significant change in occurs with the 9th zoom-in, featuring a large regular stripe, resulting in higher values, indicative of a more two-dimensional structure. The fractal dimension of the full phase space (zoom level 0) is also consistent with what we observe in Figure 3. At this scale and resolution, the majority of the regular phase space is contained within the four large islands, which lack substructures, so the fractal dimension is closer to 2. An analogous behaviour can be observed in the alternative zoom-ins of Appendix B, which targeted the border of the green regular region in the 4th zoom-in.